# Toward efficient resource utilization at edge nodes in federated learning

### v1

@@ 5 Results and Discussion
Our carefully designed experiments have yielded valuable insights: in a federated training setting with restricted resources, choosing a limited number of layers at each client location can achieve results comparable to training the complete model at each site. A noteworthy observation we made was that a negative correlation exists between the number of clients and the number of layers that need to be trained at each client location. As the number of clients increases, the required number of layers trained per client decreases, indicating that this approach can be especially effective for cross-device use cases. These findings provide valuable insights into federated learning optimization and offer practical recommendations for limited resource edge device deployment.
@@ 5.1 Experimental Settings
To explore the viability of our proposed approach, we utilized three open-source datasets and their associated machine learning models, specifically selected to represent a diverse range of application domains (computer vision, natural language processing (NLP), and Internet-of-Things). Our investigation focused on evaluating model convergence, network load and communication cost, impact on training time, and resource utilization.
@@
  • Experiment 1: This experiment focuses on a computer vision task that utilizes CIFAR-10 dataset [^2] http://www.cs.toronto.edu/~kriz/cifar.html. The dataset consists of $60,000$ colour images with dimensions of $32\times 32$ pixels, grouped into $10$ classes, with $6,000$ samples per each. The dataset is divided into two subsets: $50,000$ training images and $10,000$ test images. Also, we used the VGG16 model Krizhevsky et al. (2009) in this experiment. Table 1 shows the model architecture, including layer types, output dimensions, and the number of trainable parameters per layer. The model has a total of $14,736,714$ parameters and $14$ trainable layers, including the output layer. The generated model size is $53.5MB$. For more information, please refer to the client source code available on GitHub. [^3] https://github.com/saadiabadi/cifar_updated.git.
@@
  • Experiment 2: This experiment is centred around sentiment analysis tasks using the IMDB dataset v1.0444https://ai.stanford.edu/~amaas/data/sentiment/. The dataset consists of $50,000$ reviews, with a maximum of $30$ reviews per movie, equally divided between positive and negative reviews. In Maas et al. (2011), the dataset has been used for sentiment analysis tasks, where a machine learning model predicts whether a given review is positive or negative based on the review text. We used a deep learning model to predict the review decision for this task. Table 2 shows the detailed architecture of the model, including the layer types, dimensions and parameters used to construct and generate the initial model. Further technical details can be found on the client source code GitHub [^5] https://github.com/saadiabadi/IMDB_Example.git.
@@
  • Experiment 3: In this experiment, the focus is on human activity recognition (HAR) using the CASA dataset [^6] https://archive.ics.uci.edu/ml/datasets/Human+Activity+Recognition+from+Continuous+Ambient+Sensor+Data. The dataset comprises $13,956,534$ patterns collected over two months from 30 homes using continuous ambient and PIR sensors. Each pattern consists of a set of $37$ features linked to different sensors distributed throughout the home, representing daily human activities such as sleep, eating, reading, and watching TV. The goal of this experiment is to classify the output into $10$ different daily activities for each user. We use a Long Short-Term Memory (LSTM) model with an input layer dimension of $(100,1,36)$, four dense layers, and one output layer. The model has $68,884$ trainable parameters, and the compiled model size is $254KB$. For further technical details, please refer to the client source code available on GitHub [^7] https://github.com/saadiabadi/Casa_IoT_Example.git.
@@ 5.2.1 Model performance
Our goal is to attain accuracy similar to conventional federated learning methods while implementing a new strategy that maximizes the utilization of edge resources. This approach not only upholds data privacy, which is crucial in any federated training environment but also minimizes training costs and reduces data transfer. To demonstrate the effectiveness of this strategy, we conducted a series of experiments where we trained different numbers of randomly selected model layers. Through these experiments, we systematically increased the number of layers to assess their impact on the overall performance of the model. To compare the accuracy of centrally trained and FL global models, we varied the number of trainable layers for the FL settings. Figure 1 shows the accuracy of the VGG16 model trained on the CIFAR-10 dataset, both centrally and in FL settings (10 clients and the data split partition was equally across 10 clients), with different trainable layers randomly selected per round.
@@
The centralized model achieved an accuracy of $87.00\%$, while the FL model, with all layers included in the training process, achieved a slightly lower accuracy of $86.08\%$, a difference of only $0.92\%$. However, FL with slightly lower accuracy offers a privacy-preserving training environment.
Our experiments revealed that training $10$ randomly selected layers in each round of the model was enough to capture and learn the data behaviour, resulting around 85.70% in accuracy. Figure 1 indicates that the model began to converge from the first round of training. As we reduced the number of trained layers to approximately $50\%$ ($7$ layers) of the model, the accuracy gap increased. Nevertheless, the accuracy remained 84.79% compared to the baseline, given the reduction in cost achieved through this approach, as elaborated in the subsequent sections.
@@
To highlight the advantages of our proposed approach, we evaluated its adaptability and robustness across two distinct domains, utilizing both CASA and IMDB datasets in federated learning settings, involving 10 clients, and equal subset of data amount split from the original dataset. Figure 2 illustrates a convergence in terms of model performance for both datasets. As depicted in Figure 2-a, for the HAR task, training only $33\%$ ($2$ randomly selected layers every round) of the model resulted in good accuracy (around 79.01%) compared to the fully trained ($6$ layers, 80.20%) model with a small gap. In contrast, training $66\%$ ($4$ layers) of the model nearly achieved the same level of accuracy as training all the model layers. Interestingly, similar trends were observed in the NLP experiments, as demonstrated in Figure 2-b.
@@ 5.2.2 Trainable layer distribution
As the number of clients participating in a training round increases, the probability of engaging all layers of a model in the global training process also rises. Simultaneously, the volume of training data expands. This heightened participation and increased training data elevate the likelihood of training all model layers over each client’s data as multiple rounds progress. Consequently, this phenomenon significantly influences the model’s convergence.This effect is particularly noticeable when we randomly select different trainable layers of the model ($25\%$, $50\%$, and $75\%$) for training. During the training process, we noticed that each client has the opportunity to train every layer of the model at least once. Moreover, the distribution of layers among all clients is equitable, ensuring a balanced contribution to the training process , as demonstrated in Figure 3. The equal distribution of model layers has been used across all client’s with different training settings (i.e. $4$, $7$, and $10$ layers).
@@ 5.2.3 The impact of scaling the number of clients (edge nodes) on the model accuracy
  • Exp. 1: In this experiment, we partitioned the dataset among $10$ distinct clients to train the entire model, which consisted of $14$ trainable layers, within the context of federated learning settings.
@@
Moreover, we evaluate the effectiveness of our approach by increasing the number of clients while keeping the data amount constant to train 7 layers. As shown in Figure 6, Despite employing the same model architecture and total data size, we have observed similar performance when using either 20 or 10 clients. It is important to note that we divided the dataset into 10 and 20 partitions, which were assigned to 10 and 20 clients, respectively. By increasing the number of clients involved in constructing the global model, we noticed an improvement in the model’s performance. Therefore, our findings indicate that the global model generated by 20 clients outperformed the one built by 10 clients. Furthermore, we observed that the model trained with 20 clients achieved good accuracy, albeit slightly different from the baseline. This can be attributed to the model’s capability to learn hidden patterns from data with sufficient samples more efficiently, even when the client has a smaller sample size. These findings highlight the significance of involving more contributors (clients) and selecting an appropriate number of trainable layers in the federated learning (FL) settings, as they have a substantial impact on the model’s performance. Additionally, this outcome allows for a more precise estimation of the training budget and optimizing the utilization of available resources.
@@ 5.2.4 Training time
In the previous sections, we demonstrated the ability to maintain accuracy while freezing parts of the model. Nonetheless, it is crucial to consider the effect on training time, mainly when dealing with larger models. Nevertheless, our approach demonstrates its effectiveness in accelerating the training process. This is achieved by distributing sub-layers of the entire model among the clients to be trained over clients’ local data within the Federated Learning (FL) network, ultimately reducing training time.
Figure 7 illustrates the total training time (in minutes) required by 10 clients to complete 100 training rounds. Training the entire model takes approximately 331 minutes while training 75% of the model saves approximately 21 minutes (a 7% reduction in time). There is a significant difference of 63 minutes between training 4 layers (25% of the model) compared to training 14 layers. By training approximately 50% of the model, we can save around 36 minutes (a 10% reduction compared to training the entire model) while still maintaining accuracy (refer to Figure 1).
@@ 5.2.5 Transferred data size and number of trainable parameters
The size of transferred parameters (weights) in Federated Learning (FL) settings naturally depends on the number of trainable layers in the neural network. In addition to the benefits mentioned above, layer sub-selection has the potential to reduce the amount of resources required on the client side. This can include computing power, storage space, and network capacity needed for communication.
@@ 5.2.6 Resources utilization based on VM’s
To expedite the training process for a single client, it often demands an increase in computational resources, such as CPU and RAM. However, this can lead to higher overall training costs and limitations in scaling up these resources, potentially resulting in a single point of failure. As a solution, scaling out the number of clients with lower resources can mitigate these challenges. Training the model partially while increasing the number of clients, we’ve shown in Figure 5 that it doesn’t adversely affect the model’s accuracy. Consequently, this strategy allows for the model training process to involve constrained devices like Jetson Nano or VM’s with limited configurations (e.g., ssc.small). This approach enables the establishment of a cluster incorporating restricted nodes to train the ML model in a FL setting. The primary goal is to reduce training costs and enhance resilience against failures.
@@ 6 Conclusion
This paper introduces a novel approach for training deep learning models in a federated learning setting, aiming to efficiently utilize edge node resources and reduce network workload. The proposed approach trains a specific number of the model’s layers, randomly selected every training round, and freezes the remaining layers. This approach can enable IoT devices with restricted resources (as examplified here with the Jetson Nano) to participate in training larger models.
The approach was evaluated for three tasks: sentiment analysis using the IMDB dataset, object detection using the CIFAR-10 dataset, and human activity recognition using the CASA dataset, each with different model architectures. The experimental results demonstrate that training only a part of the model in the federated learning setting has a significant impact on resource utilization, communication, training budget, and model performance. Furthermore, increasing the number of contributors also considerably affects model performance while keeping the amount of data fixed as we demonstrated in the results.
Overall, the study demonstrates the potential of our approach to train deep learning models in a federated learning setting efficiently, enabling participation from a diverse range of devices and clients without sacrificing model performance.
@@ Acknowledgement
This work was funded by the eSSENCE strategic collaboration on eScience (Alawadi, Ait-Mlouk, Toor, and Hellander), and supported by Halmstad University. The authors also would like to thank SNIC for providing cloud resources.

### v2

@@ 4 Results and Discussion
Our carefully designed experiments have yielded valuable insights: in a federated training setting with restricted resources, choosing a limited number of layers at each client location can achieve results comparable to training the complete model at each site. A noteworthy observation we made was that a negative correlation exists between the number of clients and the number of layers that need to be trained at each client location. As the number of clients increases, the required number of layers trained per client decreases, indicating that this approach can be especially effective for cross-device use cases. These findings provide valuable insights into FL optimization and offer practical recommendations for limited resource edge device deployment.
@@ 4.1 Experimental Settings
To explore the viability of our proposed approach, we utilized three open-source datasets and their associated ML models initialized with random weight, specifically selected to represent a diverse range of application domains (computer vision, NLP, and Internet-of-Things). Our investigation focused on evaluating model convergence, network load and communication cost, impact on training time, and resource utilization.
@@
  • Experiment 1: This experiment focuses on a computer vision task that utilizes CIFAR-10 dataset [^2] http://www.cs.toronto.edu/~kriz/cifar.html. The dataset consists of $60,000$ colour images with dimensions of $32\times 32$ pixels, grouped into $10$ classes, with $6,000$ samples per each. The dataset is divided into two subsets: $50,000$ training images and $10,000$ test images. In this experiment, we randomly generated n-client data from Cifar10, ensuring that each client held an equal number of samples and that the data were independent and identically distributed (IID). Also, we used the VGG16 model Krizhevsky et al. (2009) in this experiment. Table 1 shows the model architecture, including layer types, output dimensions, and the number of trainable parameters per layer. The model has a total of $14,736,714$ parameters and $14$ trainable layers, including the output layer. The generated model size is $53.5MB$. For more information, please refer to the client source code available on GitHub. [^3] https://github.com/saadiabadi/cifar_updated.git.
@@
  • Experiment 2: This experiment is centred around sentiment analysis tasks using the IMDB dataset v1.0444https://ai.stanford.edu/~amaas/data/sentiment/. The dataset consists of $50,000$ reviews, with a maximum of $30$ reviews per movie, equally divided between positive and negative reviews. In Maas et al. (2011), the dataset has been used for sentiment analysis tasks, where a ML model predicts whether a given review is positive or negative based on the review text. We used a DL model to predict the review decision for this task. Table 2 shows the detailed architecture of the model, including the layer types, dimensions and parameters used to construct and generate the initial model. It’s important to highlight that the dataset has been randomly partitioned among n clients, where each client received the same amount of data and maintained the same class distribution, meaning all clients followed the IID distribution. Further technical details can be found on the client source code GitHub [^5] https://github.com/saadiabadi/IMDB_Example.git.
@@
  • Experiment 3: In this experiment, the focus is on human activity recognition (HAR) using the CASA dataset [^6] https://archive.ics.uci.edu/ml/datasets/Human+Activity+Recognition+from+Continuous+Ambient+Sensor+Data. The dataset comprises $13,956,534$ patterns collected over two months from 30 homes using continuous ambient and PIR sensors. Each pattern consists of a set of $37$ features linked to different sensors distributed throughout the home, representing daily human activities such as sleep, eating, reading, and watching TV Alkhabbas et al. (2020). It’s important to highlight that each home’s data represents an individual client in this setting; both the data size and the number of patterns varied among clients, resulting in a non-independent and identified distributed (Non-IID) distribution. The goal of this experiment is to classify the output into $10$ different daily activities for each user. We use a Long Short-Term Memory (LSTM) model with an input layer dimension of $(100,1,36)$, four dense layers, and one output layer. The model has $68,884$ trainable parameters distributed across $6$ trainable layers, and the compiled model size is $254KB$. For further technical details, please refer to the client source code available on GitHub [^7] https://github.com/saadiabadi/Casa_IoT_Example.git.
@@ 4.2.1 Model performance
Our goal is to attain accuracy similar to conventional FL methods while implementing a new strategy that maximizes the utilization of edge resources. This approach not only upholds data privacy, which is crucial in any federated training environment but also minimizes training costs and reduces data transfer. To demonstrate the effectiveness of this strategy, we conducted a series of experiments where we trained different numbers of randomly selected model layers. Through these experiments, we systematically increased the number of layers to assess their impact on the overall performance of the model. To compare the accuracy of centrally trained and FL global models, we varied the number of trainable layers for the FL settings. Figure 2 shows the accuracy of the VGG16 model trained on the CIFAR-10 dataset, both centrally and in FL settings (10 clients and the data split partition was equally across 10 clients), with different trainable layers randomly selected per round.
@@
The centralized model achieved an accuracy of $87.00\%$, while the FL model, with all layers included in the training process, achieved a slightly lower accuracy of $86.08\%$, a difference of only $0.92\%$. However, FL with slightly lower accuracy offers a privacy-preserving training environment. Our experiments revealed that training $10$ randomly selected layers in each round of the model was enough to capture and learn the data behaviour, resulting around 85.70% in accuracy. Figure 2 indicates that the model began to converge from the first round of training. As we reduced the number of trained layers to approximately $50\%$ ($7$ layers) of the model, the accuracy gap increased. Nevertheless, the accuracy remained 84.79% compared to the baseline, given the reduction in cost achieved through this approach, as elaborated in the subsequent sections.
@@
To highlight the advantages of our proposed approach, we evaluated its adaptability and robustness across two distinct domains, utilizing both CASA and IMDB datasets in FL settings, involving 10 clients, corresponding to 10 homes in the CASA dataset and equal subset of data amount split from the IMDB dataset. Figure 3 illustrates a convergence in terms of model performance for both datasets. As depicted in Figure 3-a, for the HAR task, training only $33\%$ ($2$ randomly selected layers every round) of the model resulted in good accuracy (around 79.01%) compared to the fully trained ($6$ layers, 80.20%) model with a small gap. In contrast, training $50\%$ ($3$ layers) or $66\%$ ($4$ layers) of the model nearly achieved the same level of accuracy as training all the model layers. Interestingly, similar trends were observed in the NLP experiments, as demonstrated in Figure 3-b.
@@ 4.2.2 Trainable layer distribution
As the number of clients participating in a training round increases, the probability of engaging all layers of a model in the global training process also rises. Simultaneously, the volume of training data expands. This heightened participation and increased training data elevate the likelihood of training all model layers over each client’s data as multiple rounds progress. Consequently, this phenomenon significantly influences the model’s convergence.This effect is particularly noticeable when we randomly select different trainable layers of the model ($25\%$, $50\%$, and $75\%$) for training. During the training process, we noticed that each client has the opportunity to train every layer of the model at least once. Moreover, the distribution of layers among all clients is equitable, ensuring a balanced contribution to the training process, as demonstrated in Figure 4. The equal distribution of model layers has been used across all clients with different training settings (i.e. $4$, $7$, and $10$ layers).
@@ 4.2.3 The impact of scaling the number of clients (edge nodes) on the model accuracy
  • Exp. 1: In this experiment, we partitioned the dataset among $10$ distinct clients to train the entire model, which consisted of $14$ trainable layers, within the context of FL settings.
@@
Moreover, we evaluate the effectiveness of our approach by increasing the number of clients while keeping the data amount constant to train 7 layers. As shown in Figure 7, Despite employing the same model architecture and total data size, we have observed similar performance when using either 20 or 10 clients. It is important to note that we divided the dataset into 10 and 20 partitions, which were assigned to 10 and 20 clients, respectively. By increasing the number of clients involved in constructing the global model, we noticed an improvement in the model’s performance. Therefore, our findings indicate that the global model generated by 20 clients outperformed the one built by 10 clients. Furthermore, we observed that the model trained with 20 clients achieved good accuracy, albeit slightly different from the baseline. This can be attributed to the model’s capability to learn hidden patterns from data with sufficient samples more efficiently, even when the client has a smaller sample size. These findings highlight the significance of involving more contributors (clients) and selecting an appropriate number of trainable layers in the FL settings, as they substantially impact the model’s performance. Additionally, this outcome allows for a more precise estimation of the training budget and optimizing the utilization of available resources.
@@ 4.2.4 Training time
In the previous sections, we demonstrated the ability to maintain accuracy while freezing parts of the model. Nonetheless, it is crucial to consider the effect on training time, mainly when dealing with larger models. Nevertheless, our approach demonstrates its effectiveness in accelerating the training process. This is achieved by distributing sub-layers of the entire model among the clients to be trained over clients’ local data within the FL network, ultimately reducing training time.
Figure 8 illustrates the total training time (in minutes) required by 10 clients to complete 100 training rounds. Training the entire model takes approximately 331 minutes while training 75% of the model saves approximately 21 minutes (a 7% reduction in time). There is a significant difference of 63 minutes between training 4 layers (25% of the model) compared to training 14 layers. By training approximately 50% of the model, we can save around 36 minutes (a 10% reduction compared to training the entire model) while still maintaining accuracy (refer to Figure 2). Take into account that the observed time reduction is relatively low, influenced by various factors such as the computational resources of individual clients, the size of the local datasets, and the presence of straggler clients, making the training round longer.
@@
Analyzing the training time across various rounds to achieve a specific accuracy level is crucial. Table 3 illustrates the time required to train 4, 7, 10 and 14 layers, respectively, and the corresponding accuracy obtained after 40, 60, 80 and 100 training rounds.
Training 4 layers selected randomly from the complete model achieves 70.54% accuracy within 40 rounds, which requires 104 minutes. Notably, increasing the number of rounds by 20 each time consistently requires around 53 minutes, demonstrating a constant trend. Furthermore, a substantial enhancement in model performance is observed, obtaining 79.02% accuracy after 100 rounds. This pattern is consistent across different layers. Variations were observed when comparing the execution time for training different layers within the same communication rounds category. For instance, the time difference between training 7 and 10 layers for 40 rounds is 6 minutes, and for 60 rounds is 9 minutes. However, the accuracy difference among different layers with the same number of rounds remained relatively close. Comparing a fully trained model (14 layers) after 100 rounds with 7 and 10 layers shows an insignificant accuracy difference, confirming that training a portion of the model while maintaining good accuracy.
Table 3: The time required (in minute) to achieve a certain level of accuracy (Acc.) during various communication rounds to train the VGG16 model by 10 clients in the FL context.
40 rounds 60 rounds 80 rounds 100 rounds
Number of Layers Time Acc. Time Acc. Time Acc. Time Acc.
4 Layers 107 70.54% 161 75.95% 214 77.72% 268 79.02%
7 Layers 118 79.95% 177 83.26% 236 84.4% 295 84.75%
10 Layers 124 81.73% 186 83.82% 248 85.08% 310 85.68%
14 Layers 132 82.79% 199 85.25% 265 85.88% 331 86.08%
@@ 4.2.5 Transferred data size and number of trainable parameters
The size of transferred parameters (weights) in the FL settings naturally depends on the number of trainable layers in the neural network. In addition to the benefits mentioned above, layer sub-selection can potentially reduce the amount of resources required on the client side. This can include computing power, storage space, and network capacity needed for communication
@@ 4.2.6 Resources utilization based on VM’s
To expedite the training process for a single client, it often demands an increase in computational resources, such as CPU and RAM. However, this can lead to higher overall training costs and limitations in scaling up these resources, potentially resulting in a single point of failure. As a solution, scaling out the number of clients with lower resources can mitigate these challenges. Training the model partially while increasing the number of clients, we’ve shown in Figure 6 that it doesn’t adversely affect the model’s accuracy. Consequently, this strategy allows for the model training process to involve constrained devices like Jetson Nano or VM’s with limited configurations (e.g., ssc.small). This approach enables the establishment of a cluster incorporating restricted nodes to train the ML model in FL setting. The primary goal is to reduce training costs and enhance resilience against failures.
@@ 4 Layers 7 Layers 10 Layers 14 Layers
4.3 Limitations of the proposed approach
The proposed approach can be implemented in any FL framework with consistent behaviour. Integrating differential privacy (DP) into our technique and comparing it with vanilla FL with DP may impact the results, but our approach maintains its fundamental characteristics. Adding DP introduces additional thresholds to both, significantly influencing training time and model accuracy.
However, it’s important to acknowledge certain limitations in the proposed approach:
  1. 1.
    The number of selected layers remains fixed during the FL training process, potentially affecting client performance and execution time in local training. This constraint could be mitigated by dynamically selecting an optimal number of layers each round based on the ideal resources for each client.
  2. 2.
    With the emergence of large language models (LLM), the approach might require additional engineering to handle the sophistication of such complex models.
  3. 3.
    Each client’s sequential training of layers represents a limitation that could be addressed through more sophisticated strategies in future iterations.
@@ 5 Conclusion
This paper introduces a novel approach for training DL models in the FL setting, aiming to efficiently utilize edge node resources and reduce network workload. The proposed approach trains a specific number of the model’s layers, randomly selected every training round, and freezes the remaining layers. This approach can enable IoT devices with restricted resources (as exemplified here with the Jetson Nano) to participate in training larger models.
The approach was evaluated for three tasks: sentiment analysis using the IMDB dataset, object detection using the CIFAR-10 dataset, and human activity recognition using the CASA dataset, each with different model architectures. The experimental results demonstrate that training only a part of the model in the FL setting has a significant impact on resource utilization, communication, training budget, and model performance. Furthermore, increasing the number of contributors also considerably affects model performance while keeping the amount of data fixed as we demonstrated in the results.
Overall, the study demonstrates the potential of our approach to train DL models in a FL setting efficiently, enabling participation from a diverse range of devices and clients without sacrificing model performance.
@@ Acknowledgement
This work was funded by the eSSENCE strategic collaboration on eScience (Alawadi, Ait-Mlouk, Toor, and Hellander), and supported by Blekinge Institute of Technology (BTH). The authors also would like to thank SNIC for providing cloud resources.